OpenStudy (anonymous):

Need help Solve the system by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4 show all your work

jigglypuff314 (jigglypuff314):

first isolate x in the third equation so 2x + 2z = 4 => 2x = -2z + 4 => x= -z + 2 then you can plug x into the first equation and reduce so -x -y -z = -8 => -(-z+2)-y-z=-8 => z-2-y-z=-8 => y=6 substitute x and y into the second equation so -4x+4y+5z=7 => -4(-z+2)+4(6)+5z=7 => 4z-8+24+5z=7 => 9z=-9 => z=-1 then you can substitute z into the third equation to find x 2x+2z=4 => 2x+2(-1)=4 => 2x-2=4 => 2x=6 => x=3 so x=3, y=6, and z=-1

OpenStudy (anonymous):

you are a life saver